# Data-oriented Dynamic Fine-tuning Parameter Selection Strategy for FISH Mask based Efficient Fine-tuning

### v1

@@ 5 Experiments Setup
Figure 2: Illustrative compared results on SST-2
@@
Table 2: Compared results on SST-2 by BERT
Method Updated Params QNLI SST-2 $\text{MNLI}_{\text{m}}$ $\text{MNLI}_{\text{mm}}$ CoLA MRPC STS-B RTE QQP
Dense Fine-tuning
100% 93.4 94.9 87.0 86.1 61.0 86.6 86.5 70.9 80.5
Bit-Fit
0.08% 90.4 94.5 85.0 84.8 60.3 86.3 85.0 69.6 78.5
FISH Mask
0.08% 93.3 94.0 85.3 84.9 56.4 86.2 85.7 70.2 79.3
Random Mask
0.50% 89.8 93.4 83.7 84.0 43.2 77.8 87.7 61.3 77.2
Diff Pruning
0.50% 91.9 93.8 86.0 85.5 61.0 86.2 85.6 67.5 80.1
FISH Mask
0.50% 93.1 94.7 86.5 85.9 61.6 87.1 86.5 71.2 80.2
IRD
0.50% 93.3 93.7 86.3 86.6 62.1 90.7 89.9 72.6 88.4
(B)FISH Mask
0.10% 89.9 91.3 79.6 80.4 60.8 88.2 88.4 71.1 85.6
(B)IRD
0.10% 90.4 91.3 82.1 82.4 61.3 88.6 88.8 71.1 86.1
(B)FISH Mask
0.02% 89.4 90.8 73.9 78.1 53.6 88 87.8 66.1 82.8
(B)IRD
0.02% 89.1 90.7 79.1 79.7 54.7 87.9 87.9 69.3 83.9
Table 3: Performance on GLUE. (B) means the foundation model is BERT-base and the default foundation model is BERT-large. Updated Params means the ratio of fine-tuning parameters to all parameters.
@@
5.1 Datasets and Baselines
Datasets. We evaluate IRD on the GLUEWang et al. (2019) benchmark dataset compared with FISH Mask method. GLUE is a multi-task benchmark that contains 10 tasks for LLMs evaluation. As task AX only has test set for zero-shot learning, we ignore it. We conduct experiments on the original data splits from GLUE.
CoLAWarstadt et al. (2018) is a 2-class classification task of English grammatical acceptability judgments.
SST-2Socher et al. (2013) is a 2-class classification task about human sentiment in movie reviews.
MRPCDolan and Brockett (2005) is a 2-class classification task about senteces pair are semantically equivalent or not.
STS-BCer et al. (2017) is a regression task about sentences pair similarity score from 1.0 to 5.0.
MNLIWilliams et al. (2018) is a 3-class classification task about the relationship between premise and hypothesis sentences.
QNLIRajpurkar et al. (2016) is a 2-class classification task about the relationship between question and answer.
WNLILevesque et al. (2012) is a 2-class classification task for predicting whether the original sentence entails the sentence with the pronoun substituted. This task only contains 635 samples for training.
Baselines. Since IRD is based on the selective PEFT method. We choose some selective methods as baselines, including Bit-FitZaken et al. (2022), Diff PruningGuo et al. (2021), original FISH MaskSung et al. (2021). Also, following the settings of Sung et al. (2021), we introduce some simple settings as baselines, such as Dense Fine-tuning and Random Mask. Because the data set and experimental settings are exactly the same, we quoted some experimental results in Sung et al. (2021) and focused on testing IRD and the original FISH Mask method.
5.2 Evaluation Metrics
Evaluation Metrics. The metrics of CoLA task is Matthews correlation coefficientMatthews (1975). The metrics of STS-B task is Pearson and Spearman correlation coefficients. The metrics of MRPC and QQP tasks are combined score(half the sum of the $F_{1}$ and Accuracy). The metrics of other tasks are Accuracy. The settings of evaluation metrics is following Sung et al. (2021).
@@ 5.3 Implementation Details
We only use the validation set results as the final results because of the limitation of 2-time uploads to the GLUE official website each day. We follow the same parameter settings for the classification task as BERTDevlin et al. (2019). Apart from the BERT model, we also compare IRD with FISH Mask in the auto-regressive model like GPT-2Radford et al. (2019). We use early stopping to prevent overfit and underfit. The prompt template for GPT-2 is following the Table 1 in Zhong et al. (2023) .
@@
Task max_seq_length
SST-2 128
CoLA 128
RTE 384
STS-B 256
WNLI 128
MRPC 128
QQP 512
QNLI 512
MNLI 512
@@
For the early stopping mechanism, we uses a patience parameter of 10 epochs and a threshold value of 0.3 to determine stopping criteria. Training is halted if there’s no improvement exceeding the patience parameter over 10 consecutive epochs and if the metric value exceeds the threshold value of 0.3. The $learning\_rate$ is 5e-5 in most of the tasks, except WNLI is 5e-6. If the training context only has one sentence, we use "[CLS] sentence I". Otherwise we use "[CLS] sentence I [SEP] sentence II" to concatenate each context. The $batch\_size$ is set to 32. The loss function is cross-entropy and the optimizer is AdamKingma and Ba (2015).
@@ 5.4.2 Results with BERT on GLUE
Table 3 shows the detailed performance of IRD and baseline methods on the GLUE benchmark. We conducted seven experiments based on BERT-large foundation model. IRD achieves the best performance on 7 out of 9 GLUE subtasks. Besides, under the BERT-base foundation model, IRD achieves the best performance on 6 out of 9 GLUE subtasks.
@@ 5.4.3 Results with GPT-2 on GLUE
Figure 5: Contrastive Study
@@ 5.4.4 Contrastive Study
To demonstrate the enhancement of IRD over the FISH Mask method, we switch the sample choosing step (Step 4) and parameter choosing step(Step 7) of Algorithm1 to a contrastive way, aiming for inverse results. In Algorithm1, Step 4 represents selecting the larger half of elements from set $X_{l}$ to form a new set. Here, we modify it to select the larger half of elements from set $X_{l}$, but this time from the latter half (Shown in Eq.5). In Eq.6, we make the same operation for the set of $\Theta_{l}^{*}$.
$X_{l+1}=\{x_{i}\in X_{l}|\hat{F}_{x_{i}}<{\rm sort}(\hat{F}_{x_{i}})_{\frac{|X_{l}|}{2}}\}$ (5)
$\Theta_{l+1}^{*}=\{\theta_{i}\in\Theta_{l}|\hat{F}_{\theta_{i}}<{\rm sort}(\hat{F}_{\theta})_{\frac{|\Theta_{l}|}{2}}\}$ (6)
@@
In non-tie results, IRD is better than Fish-mask. We can see the obvious advantages of IRD from Fig.4. We set IRD to search in the opposite direction of optimization, that is, search for worse values in each iteration, and put the results on the right, and IRD on the left for comparison in Fig.4. Experimental results prove that the IRD method is better than reverse IRD on four subtasks. There are far more downward arrows than upward arrows (20>12, indicating that reverse IRD reduces performance), which proves the reasonableness and superiority of IRD search.
@@ 6 Conclusion
Limitations
Although we proposed a general data-oriented PEFT method optimization idea, the algorithm we designed only optimizes a specific model. Furthermore, IRD cannot dynamically adjust the model during training, and all settings are determined before training. Both of these limitations can be further optimized in subsequent work.
Ethics Statement
In this study on parameter efficient fine tuning using the GLUE benchmark and pre-trained models like BERT or GPT-2, we recognize the importance of ethical considerations in our methodology and findings. Acknowledging that the datasets and models employed may inherently carry biases reflective of their training sources, we are mindful of the potential for misuse or misinterpretation in sensitive applications. We advocate for the responsible application of our findings, emphasizing the importance of ethical usage to avoid social harm.
We have ensured the transparency and reproducibility of our research by making our code and methodologies publicly available. This is in line with our commitment to fostering trust and integrity in natural language processing research.
Acknowledgments
Appendix A Tables of Precision Results
FISH Mask on CoLA
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.5364-
12896
0.1 0.6083- 0.5942-
70776 58898
0.5 0.6196- 0.6137-
50422 21879
2.5 0.6211- 0.6120-
92519 85127
IRD on CoLA
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.5469-
58705 $\uparrow$
0.1 0.6132- 0.6132-
37959 $\uparrow$ 37959 $\uparrow$
0.5 0.6194- 0.6240-
38412 $\downarrow$ 83897 $\uparrow$
2.5 0.6162- 0.6120-
36749 $\downarrow$ 85127
Table 5: Compared results on CoLA by BERT
FISH Mask on RTE
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.6606-
49819
0.1 0.7111- 0.7400-
91336 72202
0.5 0.7364- 0.7364-
62094 62094
2.5 0.7148- 0.7148-
01444 01444
IRD on RTE
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.6931-
40794 $\uparrow$
0.1 0.7111- 0.7111-
91336 91336 $\downarrow$
0.5 0.7509- 0.7400-
02527 $\uparrow$ 72202 $\uparrow$
2.5 0.7328- 0.7148-
51986 $\uparrow$ 01444
Table 6: Compared results on RTE by BERT
FISH Mask on STS-B
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.8784-
29404
0.1 0.8843- 0.8874-
99414 82622
0.5 0.8904- 0.8899-
70230 26925
2.5 0.8930- 0.8966-
60770 27263
IRD on STS-B
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.8788-
92200 $\uparrow$
0.1 0.8876- 0.8875-
48950 $\uparrow$ 52835 $\uparrow$
0.5 0.8920- 0.8921-
09119 $\uparrow$ 15572 $\uparrow$
2.5 0.8968- 0.8966-
80132 $\uparrow$ 27263
Table 7: Compared results on STS-B by BERT
FISH Mask on WNLI
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.5774-
64789
0.1 0.5633- 0.5774-
80282 64789
0.5 0.5915- 0.5915-
49296 49296
2.5 0.5211- 0.5070-
26761 42254
IRD on WNLI
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.5633-
80282 $\downarrow$
0.1 0.5633- 0.5633-
80282 80282 $\downarrow$
0.5 0.5633- 0.5633-
80282 $\downarrow$ 80282 $\downarrow$
2.5 0.5070- 0.5070-
42254 $\downarrow$ 42254
Table 8: Compared results on WNLI by BERT
FISH Mask on MRPC
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.8800-
59185
0.1 0.8817- 0.8950-
27159 19196
0.5 0.8869- 0.8825-
48529 92061
2.5 0.8747- 0.8866-
46450 20732
IRD on MRPC
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.8787-
48558 $\downarrow$
0.1 0.8862- 0.8864-
88318 $\uparrow$ 55108 $\downarrow$
0.5 0.8874- 0.8834-
31777 $\uparrow$ 27344 $\uparrow$
2.5 0.8785- 0.8866-
74346 $\uparrow$ 20732
Table 9: Compared results on MRPC by BERT
FISH Mask on QQP
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.8279-
72035
0.1 0.8562- 0.8630-
98482 70508
0.5 0.8766- 0.8771-
64736 79970
2.5 0.8847- 0.8873-
38135 17030
IRD on QQP
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.8386-
33700 $\uparrow$
0.1 0.8616- 0.8616-
14642 $\uparrow$ 14642 $\downarrow$
0.5 0.8820- 0.8768-
65026 $\uparrow$ 32596 $\downarrow$
2.5 0.8845- 0.8873-
00562 $\downarrow$ 17030
Table 10: Compared results on QQP by BERT
FISH Mask on QNLI
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.8943-
80377
0.1 0.8993- 0.9059-
22716 12502
0.5 0.9081- 0.9062-
09098 78602
2.5 0.9086- 0.9084-
58246 75197
IRD on QNLI
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.8909-
02435 $\downarrow$
0.1 0.9048- 0.9040-
14205 $\uparrow$ 82006 $\downarrow$
0.5 0.9070- 0.9075-
10800 $\downarrow$ 59949 $\uparrow$
2.5 0.9117- 0.9084-
70090 $\uparrow$ 75197
Table 11: Compared results on QNLI by BERT
FISH Mask on MNLI-m
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.7388-
69078
0.1 0.7963- 0.8159-
32145 95925
0.5 0.8295- 0.8300-
46612 56037
2.5 0.8382- 0.8347-
06826 42741
IRD on MNLI-m
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.7909-
32247 $\uparrow$
0.1 0.8207- 0.8207-
84513 $\uparrow$ 84513 $\uparrow$
0.5 0.8268- 0.8255-
97606 $\downarrow$ 73102 $\downarrow$
2.5 0.8368- 0.8347-
82323 $\downarrow$ 42741
Table 12: Compared results on MNLI-m by BERT
FISH Mask on MNLI-mm
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.7506-
10252
0.1 0.8041- 0.8189-
09032 58503
0.5 0.8251- 0.8341-
62734 13100
2.5 0.8347- 0.8359-
23352 43857
IRD on MNLI-mm
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.7967-
86005 $\uparrow$
0.1 0.8240- 0.8197-
43938 $\uparrow$ 72172 $\uparrow$
0.5 0.8227- 0.8215-
21725 $\downarrow$ 01221 $\downarrow$
2.5 0.8327- 0.8359-
90887 $\downarrow$ 43857
Table 13: Compared results on MNLI-mm by BERT
FISH Mask on SST-2
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.9071-
10092
0.1 0.9139- 0.9151-
90826 37615
0.5 0.9139- 0.9139-
90826 90826
2.5 0.9139- 0.9139-
90826 90826
IRD on SST-2
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.9071-
10092
0.1 0.9082- 0.9094-
56881 $\downarrow$ 03670 $\downarrow$
0.5 0.9162- 0.9162-
84404 $\uparrow$ 84404 $\uparrow$
2.5 0.9128- 0.9139-
44037 $\downarrow$ 90826
Table 14: Compared results on SST-2 by GPT-2
FISH Mask on CoLA
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.4411-
21566
0.1 0.4608- 0.5203-
37120 12159
0.5 0.5469- 0.5656-
07507 29691
2.5 0.5422- 0.5176-
27824 61792
IRD on CoLA
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.4471-
42466 $\uparrow$
0.1 0.5111- 0.5111-
15462 $\uparrow$ 15462 $\downarrow$
0.5 0.5282- 0.5353-
40425 $\downarrow$ 92581 $\downarrow$
2.5 0.5294- 0.5176-
39529 $\downarrow$ 61792
Table 15: Compared results on CoLA by GPT-2
FISH Mask on RTE
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.6859-
20578
0.1 0.6750- 0.6931-
90253 40794
0.5 0.6895- 0.6750-
30686 90253
2.5 0.6787- 0.6606-
00361 49819
IRD on RTE
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.6859-
20578
0.1 0.6895- 0.6895-
30686 $\uparrow$ 30686 $\downarrow$
0.5 0.6859- 0.6678-
20578 $\downarrow$ 70036 $\downarrow$
2.5 0.6642- 0.6606-
59928 $\downarrow$ 49819
Table 16: Compared results on RTE by GPT-2
FISH Mask on STS-B
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.8628-
15696
0.1 0.8726- 0.8733-
16696 94962
0.5 0.8707- 0.8714-
61240 91968
2.5 0.8741- 0.8778-
62863 00614
IRD on STS-B
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.8638-
74896 $\uparrow$
0.1 0.8727- 0.8727-
16163 $\uparrow$ 13879 $\downarrow$
0.5 0.8721- 0.8721-
18050 $\uparrow$ 98328 $\uparrow$
2.5 0.8774- 0.8778-
09002 $\uparrow$ 00614
Table 17: Compared results on STS-B by GPT-2
FISH Mask on WNLI
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.5633-
80282
0.1 0.6056- 0.5774-
33803 64789
0.5 0.6056- 0.6056-
33803 33803
2.5 0.6056- 0.6197-
33803 18310
IRD on WNLI
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.5915-
49296 $\uparrow$
0.1 0.5915- 0.5915-
49296 $\downarrow$ 49296 $\uparrow$
0.5 0.6056- 0.6056-
33803 33803
2.5 0.6197- 0.6197-
18310 $\uparrow$ 18310
Table 18: Compared results on WNLI by GPT-2
FISH Mask on MRPC
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.8308-
04639
0.1 0.8387- 0.8554-
55199 90854
0.5 0.8487- 0.8484-
79486 58462
2.5 0.8647- 0.8580-
87582 36850
IRD on MRPC
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.8226-
50646 $\downarrow$
0.1 0.8556- 0.8556-
49149 $\uparrow$ 49149 $\uparrow$
0.5 0.8498- 0.8487-
07103 $\uparrow$ 79486 $\uparrow$
2.5 0.8741- 0.8580-
56321 $\uparrow$ 36850
Table 19: Compared results on MRPC by GPT-2
IRD on SST-2
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.9116-
97248
0.1 0.9105- 0.9094-
50459 03670
0.5 0.9162- 0.9162-
84404 84404
2.5 0.9162- 0.9197-
84404 24771
IRD-inverse on SST-2
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.9013-
76147 $\downarrow$
0.1 0.9071- 0.9071-
10092 $\downarrow$ 10092 $\downarrow$
0.5 0.9151- 0.9151-
37615 $\downarrow$ 37615 $\downarrow$
2.5 0.9174- 0.9197-
31193 $\uparrow$ 24771
Table 20: Contrastive Study on SST-2 by GPT-2
IRD on CoLA
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.4260-
24490
0.1 0.5241- 0.5241-
75362 75362
0.5 0.5235- 0.5258-
22165 66331
2.5 0.5312- 0.5365-
01720 00716
IRD-inverse on CoLA
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.4440-
81207 $\uparrow$
0.1 0.4758- 0.4758-
82812 $\downarrow$ 82812 $\downarrow$
0.5 0.5206- 0.5206-
02684 $\downarrow$ 02684 $\downarrow$
2.5 0.5312- 0.5365-
51877 $\uparrow$ 00716
Table 21: Contrastive Study on CoLA by GPT-2
IRD on RTE
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.6787-
00361
0.1 0.6750- 0.6750-
90253 90253
0.5 0.6895- 0.6859-
30686 20578
2.5 0.6750- 0.6714-
90253 80144
IRD-inverse on RTE
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.6859-
20578 $\uparrow$
0.1 0.6823- 0.6823-
10469 $\uparrow$ 10469 $\uparrow$
0.5 0.6859- 0.6859-
20578 $\downarrow$ 20578
2.5 0.6823- 0.6714-
10469 $\uparrow$ 80144
Table 22: Contrastive Study on RTE by GPT-2
IRD on STS-B
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.8634-
49302
0.1 0.8677- 0.8678-
95376 00566
0.5 0.8679- 0.8680-
89910 01243
2.5 0.8706- 0.8702-
74675 57955
IRD-inverse on STS-B
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.8611-
47175 $\downarrow$
0.1 0.8684- 0.8684-
53074 $\uparrow$ 55518 $\uparrow$
0.5 0.8695- 0.8693-
07693 $\uparrow$ 85954 $\uparrow$
2.5 0.8689- 0.8702-
25294 $\downarrow$ 57955
Table 23: Contrastive Study on STS-B by GPT-2
IRD on WNLI
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.6056-
33803
0.1 0.6056- 0.6056-
33803 33803
0.5 0.6056- 0.6056-
33803 33803
2.5 0.5774- 0.5774-
64789 64789
IRD-inverse on WNLI
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.5915-
49296 $\downarrow$
0.1 0.5915- 0.5915-
49296 $\downarrow$ 49296 $\downarrow$
0.5 0.6056- 0.6056-
33803 33803
2.5 0.5774- 0.5774-
64789 64789
Table 24: Contrastive Study on WNLI by GPT-2
IRD on MRPC
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.8478-
51931
0.1 0.8389- 0.8389-
75523 75523
0.5 0.8510- 0.8510-
35571 35571
2.5 0.8531- 0.8594-
32784 27609
IRD-inverse on MRPC
Mask FISH Mask samples
Sparsity 1 16 32 128
0.02 0.8404-
77338 $\downarrow$
0.1 0.8374- 0.8374-
62481 $\downarrow$ 62481 $\downarrow$
0.5 0.8498- 0.8498-
07103 $\downarrow$ 07103 $\downarrow$
2.5 0.8651- 0.8594-
64869 $\uparrow$ 27609
Table 25: Contrastive Study on MRPC by GPT-2

### v2

@@ 5 Experiments Setup
Method Params Ratio QNLI SST-2 $\text{MNLI}_{\text{m}}$ $\text{MNLI}_{\text{mm}}$ CoLA MRPC STS-B RTE QQP
Dense Fine-tuning 100% 93.4 94.9 87.0 86.1 61.0 86.6 86.5 70.9 80.5
Bit-Fit 0.08% 90.4 94.5 85.0 84.8 60.3 86.3 85.0 69.6 78.5
FISH Mask 0.08% 93.3 94.0 85.3 84.9 56.4 86.2 85.7 70.2 79.3
Random Mask 0.50% 89.8 93.4 83.7 84.0 43.2 77.8 87.7 61.3 77.2
Diff Pruning 0.50% 91.9 93.8 86.0 85.5 61.0 86.2 85.6 67.5 80.1
LoRA 0.50% 93.5 94.0 86.7 86.4 65.5 91.4 89.7 76.9 89.4
FISH Mask 0.50% 93.1 94.7 86.5 85.9 61.6 87.1 86.5 71.2 80.2
IRD 0.50% 93.3 93.7 86.3 86.6 62.1 90.7 89.9 72.6 88.4
(B)LoRA 0.10% 90.2 91.1 81.7 82.3 59.3 88.1 84.8 67.5 86.9
(B)FISH Mask 0.10% 89.9 91.3 79.6 80.4 60.8 88.2 88.4 71.1 85.6
(B)IRD 0.10% 90.4 91.3 82.1 82.4 61.3 88.6 88.8 71.1 86.1
(B)LoRA 0.04% 88.8 91.6 79.7 80.0 54.6 87.9 84.5 61.0 85.2
(B)LoRA 0.02% 90.6 89.9 77.8 78.8 50.8 90.1 83.9 66.8 83.7
(B)FISH Mask 0.02% 89.4 90.8 73.9 78.1 53.6 88 87.8 66.1 82.8
(B)IRD 0.02% 89.1 90.7 79.1 79.7 54.7 87.9 87.9 69.3 83.9
Table 2: Performance on GLUE. (B) means the foundation model is BERT-base and the default foundation model is BERT-large. Updated Params means the ratio of fine-tuning parameters to all parameters. The Bold numbers make out the best result in the comparison with same fine-tuning ratio of parameters.
5.1 Datasets and Baselines
Datasets. We evaluate IRD on the GLUE [33] benchmark dataset compared with FISH Mask method. GLUE is a multi-task benchmark that contains 10 tasks for LLMs evaluation. As task AX only has test set for zero-shot learning, we ignore it. We conduct experiments on the original data splits from GLUE.
CoLA [35] is a 2-class classification task of English grammatical acceptability judgments.
SST-2 [30] is a 2-class classification task about human sentiment in movie reviews.
MRPC [9] is a 2-class classification task about senteces pair are semantically equivalent or not.
STS-B [4] is a regression task about sentences pair similarity score from 1.0 to 5.0.
MNLI [36] is a 3-class classification task about the relationship between premise and hypothesis sentences.
QNLI [28] is a 2-class classification task about the relationship between question and answer.
WNLI [17] is a 2-class classification task for predicting whether the original sentence entails the sentence with the pronoun substituted. This task only contains 635 samples for training.
Baselines. Since IRD is based on the selective PEFT method. We choose some selective methods as baselines, including Bit-Fit [38], Diff Pruning [12], original FISH Mask [31]. Also, following the settings of Sung et al. [31], we introduce some simple settings as baselines, such as Dense Fine-tuning and Random Mask. Because the data set and experimental settings are exactly the same, we quoted some experimental results in Sung et al. [31] and focused on testing IRD and the original FISH Mask method. In addition, we selected the widely used LoRA [14] method as our comparison benchmark to illustrate the effectiveness of our proposed method.
5.2 Evaluation Metrics
Evaluation Metrics. The metrics of CoLA task is Matthews correlation coefficient [23]. The metrics of STS-B task is Pearson and Spearman correlation coefficients. The metrics of MRPC and QQP tasks are combined score(half the sum of the $F_{1}$ and Accuracy). The metrics of other tasks are Accuracy. The settings of evaluation metrics is following Sung et al. [31].
Figure 2: llustrative compared results on SST-2
@@ 5.2 Evaluation Metrics
Table 3: Accurate compared results on SST-2 by BERT (In order to display more accurate values, the numbers in the table will be wrapped.)
@@ 5.3 Implementation Details
We only use the validation set results as the final results because of the limitation of 2-time uploads to the GLUE official website each day. We follow the same parameter settings for the classification task as BERT [6]. Apart from the BERT model, we also compare IRD with FISH Mask in the auto-regressive model like GPT-2 [27]. We use early stopping to prevent overfit and underfit. The prompt template for GPT-2 is following the Table 1 in Zhong et al. [40] .
@@
Task max_seq_length Task max_seq_length Task max_seq_length
SST-2 128 CoLA 128 RTE 384
STS-B 256 WNLI 128 MRPC 128
QQP 512 QNLI 512 MNLI 512
@@
For the early stopping mechanism, we uses a patience parameter of 10 epochs and a threshold value of 0.3 to determine stopping criteria. Training is halted if there’s no improvement exceeding the patience parameter over 10 consecutive epochs and if the metric value exceeds the threshold value of 0.3. The $learning\_rate$ is 5e-5 in most of the tasks, except WNLI is 5e-6. If the training context only has one sentence, we use "[CLS] sentence I". Otherwise we use "[CLS] sentence I [SEP] sentence II" to concatenate each context. The $batch\_size$ is set to 32. The loss function is cross-entropy and the optimizer is Adam [16].
Figure 5: Contrastive Study
@@ 5.4.2 Results with BERT on GLUE
Table 2 shows the detailed performance of IRD and baseline methods on the GLUE benchmark. We conducted 8 experiments based on BERT-large foundation model. With the scale of 0.5% parameters, LoRA achieves the best performance on 6 out of 9 GLUE subtasks. Besides, under the BERT-base foundation model with 0.1% parameters to be fine-tuned, IRD achieves the best performance on 8 out of 9 GLUE subtasks. On a smaller parameter scale (0.02%), IRD can achieve similar performance to LoRA with only half the amount of fine-tuning parameters, and is better than FISH Mask. The parameter scale of LoRA is affected by rank and cannot be linearly adjusted to 0.02%.
@@ 5.4.3 Results with GPT-2 on GLUE
Figure 6: Results with LLaMA
5.4.4 Results with LLaMA on GLUE
To further investigate the efficacy of IRD on LLaMA3.2, we conduct experiments using the LLaMA pre-trained model. Fig. 6 presents the results on the GLUE benchmark. Similar to the GPT-2 experiments, we adjust the $max_{s}eq_{l}ength$ parameter for each task. Due to computational constraints, we focus on four tasks from the GLUE benchmark. On the MRTC, COLA, and STS-B tasks, IRD demonstrates superior performance. This suggests that the benefits of IRD might be more pronounced in larger models like LLaMA. This further supports the effectiveness and generalizability of the IRD strategy across different foundation model architectures and scales. We used the full parameter search of the LLaMA 1B model, and the experimental results verified the effectiveness of the IRD method. On larger LLMs, we can use IRD for several deeper layers of the model to improve efficiency. Many existing studies have also confirmed that the last few layers of the model contain more common sense knowledge information. [34]
@@ 5.4.5 Contrastive Study
To demonstrate the enhancement of IRD over the FISH Mask method, we switch the sample choosing step (Step 4) and parameter choosing step(Step 7) of Algorithm 1 to a contrastive way, aiming for reverse results. In Algorithm 1, Step 4 represents selecting the larger half of elements from set $X_{l}$ to form a new set. Here, we modify it to select the larger half of elements from set $X_{l}$, but this time from the latter half (Shown in Eq. 5). In Eq. 6, we make the same operation for the set of $\Theta_{l}^{*}$.
$X_{l+1}=\{x_{i}\in X_{l}|\hat{\mathcal{I}}_{x_{i}}<{\rm sort}(\hat{\mathcal{I}}_{x_{i}})_{\frac{|X_{l}|}{2}}\}$ (5)
$\Theta_{l+1}^{*}=\{\theta_{i}\in\Theta_{l}|\hat{\mathcal{I}}_{\theta_{i}}<{\rm sort}(\hat{\mathcal{I}}_{\theta})_{\frac{|\Theta_{l}|}{2}}\}$ (6)
@@
In non-tie results, IRD is better than FISH mask. We can see the obvious advantages of IRD from Fig.4. We set IRD to search in the opposite direction of optimization, that is, search for worse values in each iteration, and put the results on the right, and IRD on the left for comparison in Fig.4. Experimental results prove that the IRD method is better than reverse IRD on four subtasks. There are far more downward arrows than upward arrows (20>12, indicating that reverse IRD reduces performance), which proves the reasonableness and superiority of IRD search.
@@
5.5 Compared with LoRA
As a PEFT method, LoRA[14] has attracted a huge amount of attention for its low computing cost and satisfactory performance. LoRA and methods based on it [5] became a widely compared benchmark. In order to verify the effectiveness of the IRD algorithm, we compared it with LoRA on a variety of different parameter scales. The experimental results are shown in Table 2. It shows that LoRA achieves better performance when using 0.5% parameter of BERT. Meanwhile, IRD shows better performance when using 0.1% parameter of foundation model. Since the parameter scale of LoRA depends on the rank of additive matrix, we set the rank as 1 and get the least parameter scale of LoRA on BERT base without changing the other settings (number of layers or number of matrix). LoRA only achieved similar results to IRD with twice the parameter scale (0.04% versus 0.02%). It reveals that the fine-tuning effect is not as good as IRD when the parameter scale is small.
Based on the structure of the LoRA model, it has two disadvantages compared to IRD: (1) Limited by the rank attribute of LoRA, it is impossible to finely adjust the parameter scale linearly. (2) The fine-tuning performance at a smaller parameter scale is not as good as that of IRD. (3) Compared with the selective-based PEFT method, LoRA requires additional parameter storage space. At the same time, we also believe that LoRA does have better computing efficiency and lower computational complexity.
@@ 6 Conclusion
7 Acknowledgement
This work is partially supported by Hubei Provincial Natural Science Foundation (No.2023AFB487), China Postdoctoral Science Foundation (No.2023M731253), the Key Research and Development Program of Hubei Province (2020BAB017)
[…]